# Range-Limited Heaps’ Law for Functional DNA Words in the Human Genome

### v1

@@ 5 Artificial units generated by sampling tokens randomly and a systematic deviation from the power-law
The random sampling of tokens can also be carried out on per chromosome basis. Fig.5 shows the type-token plot for 23 individual chromosomes. Protein domains observed only within one chromosome are pooled, then randomly sampled without replacement, and distinct types are counted. Chromosome 19 again stands out as the chromosome with the lowest diversity and lesser number of types. The plateaus reached for each chromosome indicate that the dictionary size has been reached. This finite size effect is different from the slight convexity of the type-token curve.
@@
The lines/curves for individual chromosomes in Fig.5 behave similarly in terms of slope before reaching the plateau. The y-intercept is very different for chromosome 19, consistent with Fig.1 and Fig.2, while those of other chromosomes are comparable. The shape of per-chromosome type-token scatter plots obtained by random sampling, which is roughly a power-law with a slight concave curving followed by a plateau, is a typical example of finite size effect. Due to the nature of generated data, the number of token is not limited whereas the number of types is finite, leading to plateaus.
@@ 6 Discussion
Since the genome size is limited, the total number of protein-domain-coding loci is limited, there is an upper limit in the #token axis. Similarly, it is not practical to use short genome regions to count #token and #type as half of the human genome are in non-coding intergenic regions. This leads to a lower limit in #token axis. Checking type-token relation in the human genome has to be range limited. On the other hand, our simulation in Fig.4 shows that beyond certain range, one should not expect power-law relationship between type and token anymore, and Heaps’ law will break down.
@@
Studies along the same lines with the present investigation include (Nasir et al., 2017; Caetano-Anollés, 2021). These works focus on phylogenetics of different species, and the word used is protein fold superfamilies (FSF). FSF information was provided by the SCOP (structural classification of proteins) database (Murzin et al., 1995; Andreeva et al., 2020). Even if SCOP domain superfamilies and Pfam domains are somewhat similar, the sample points in a type-token plot in (Nasir et al., 2017; Caetano-Anollés, 2021) are different from those in the current study. In (Nasir et al., 2017; Caetano-Anollés, 2021), each point represents a species whereas in our study a point is a chromosome or chromosome arm. When all species are represented in one plot in (Nasir et al., 2017; Caetano-Anollés, 2021), including viruses, archaea, bacteria, eukarya, the scattering of the points (genomes) in log-log scale does not follow one single straight line, but a curve.
@@
We have confirmed the result of some previous works that if tokens are randomly sampled from a dictionary (type pool), while type frequency (token per type) follows the Zifp’s law or some modification of the Zipf’s law, the resulting type-token plots, though may have a look of power-law/Heaps’ law, are actually curved (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). By simulation, we also numerically determined the impact of finite dictionary size, by removing the rare CRCPD types from consideration: the type-token plot becomes more curved with a reduction of dictionary size. The previous attempts to derive the type-token relation, with finite or infinite dictionary size do not present the result in closed-form analytic expression (Eliazar, 2011; Lü et al., 2013; Tunnicliffe and Hunter, 2022). Our quadratic regression provides a simple formula for an empirical description of the systematic deviation from the Heaps’ law. Adding a second term to better fit the result is a common practice in data analysis (Frappat et al., 2003; Li et al., 2010).
@@
The definition of “type” or “dictionary word”, or in our case, CRCPD, is supposed to be given by the database, and synonym is not merged. For example, there are 13 protein domain names containing the prefix of zf-C2H2, including the top ranking zf-C2H2 (5493 appearances in the genome) and second ranking zf-C2H2_6 (1609 appearance in the genome). The reason of not combining potential synonyms is that it should be a feature of a “dictionary”, and actually, Heaps’ law partially reflects this lexical diversity. If one would combine CRCPD that are “similar” (up to certain degree), the type-token plot and Heaps’ law in human languages should also be modified accordingly.
@@
As argued in (Lynch and Conery, 2003), a consequence of smaller population size is that selection pressure is much less effective in removing individuals without a perfect fit for the given environment. As nonadaptive forces dominated, genomes became less compact, with more redundancies, and inefficiencies were tolerated. On the other hand, this near-neutral random walk in Wright’s landscape allows a species to reach a new adaptive peak, potentially higher than the previous one (Wagner, 2019). Microorganisms have populations several orders of magnitude larger than animals. The higher level of genome redundancy and complexity in animals, versus the lower level of genome complexity in microorganisms, is then be expected. For a genome without redundancy, the number of token is roughly equal to the number of types, and Heaps exponent is one. With redundancy, exponent is expected to be less than 1.
It is suggested in (Caetano-Anollés et al., 2017; Caetano-Anollés, 2021) that viral, archaeal and bacterial proteomics have a higher level of vocabulary growth than eukarya, as indicated by the higher value of $\alpha$ exponents (around 0.8). The $\alpha$ exponent for eukarya genomes in (Caetano-Anollés, 2021) is only around 0.1-0.3, in contrast with our exponent $\alpha\sim 0.7$ within one genome. Because the higher the amount of duplication and larger the resulting gene families, the less the type-to-token ratio, we expect genomes like the human genome with abundance of duplication to have a lower $\alpha$ exponent than genomes lacking duplication. In our data and result, human, mouse, and zebrafish exhibit a lower value of $\alpha$ than those for chicken and drosophila. This lower $\alpha$ exponent is indicative of a larger amount of duplicated protein domains which is typical in presumably higher organisms.
@@
Despite the deviation from a perfect power-law for type-token plot, or in an alternative wording, that the power-law Heaps’ law is only technically correct in a limited range, there are other important lessons in this study from the linguistic perspective. Just like human languages with hundreds of new English words being created annually, albeit some of them are synonyms to the existing words, new protein domains have been generated during evolution. Even if the new protein domains may not be completely new, being at various degress similar to the old ones, these may not be that different from the new synonymous words in human languages. Some of the seemingly non-novel new protein domains may continue to evolve and eventually lead to a new meaning and a new structure.

### v2

@@ 4.2 Random sampling from CRCPD pool and a systematic deviation from the power-law
The random sampling of tokens can also be carried out on per chromosome basis. Fig.5 shows the type-token plot for 23 individual chromosomes. Protein domains observed only within one chromosome are pooled, then randomly sampled without replacement, and distinct types are counted. Chromosome 19 again stands out as the chromosome with the lowest diversity and lesser number of types. The plateaus reached for each chromosome indicate that the dictionary size has been reached. This finite size effect is different from the slight convexity of the type-token curve in Fig.4.
@@
The lines/curves for individual chromosomes in Fig.5 behave similarly in terms of slope before reaching the plateau. The y-intercept is very different for chromosome 19, consistent with Fig.1 and Fig.2, while those of other chromosomes are comparable. The shape of per-chromosome type-token scatter plots obtained by random sampling, which is roughly a power-law with a slight concave curving followed by a plateau, is a typical example of finite size effect. Due to the nature of generated data, the number of tokens is not limited whereas the number of types is finite, leading to plateaus.
Figure 4: (A) Token(x-axis)-type(y-axis) plot for randomly generated token from the whole token pool of 62299 CRCPDs. The plot is fitted by both linear regression (after logarithm transformation of both axes) and quadratic regression. Quadratic regression visibly fits the data better than linear regression, indicating a deviation from the Heaps’ law. The data points from Fig.1(C) are redrawn here (lightblue) for comparison (for both chromosomes and chromosome arms, except for Chr19). (B) Repeating the same random sampling of token and quadratic regression fitting, when the dictionary size is gradually reduced. The y-axis shows the number of CRCPD type (from top to bottom): original data; half of the singleton CRCPD types (those that appear in the human genome only once) are removed; all singletons are removed; all singletons and half of the doublet CRCPD types are removed; all singletons and all doublets are removed. The x-axis is the regression coefficient for the quadratic term.
Figure 5: Type-token plots for randomly sampled CRCPDs in individual chromosomes. Note that Chr19 (lower purple curve) distinguishes itself from the rest (see discussion in the text).
5 Results for other animal genomes
5.1 Sampling from chromosomal units
Figure 6 shows the type-token scatter plot in log-log scale with points both representing a chromosome (black) or half of a chromosome (blue). The practice of using half chromosomes is to increase the sample size and to increase the covering range of the x-axis.
To summarize, the results from these genomes where the protein domain track is available, Heaps’ law is generally true with exponent $\alpha<1$, despite potentially data quality issues such as the UniProt domain information, limited range of x-axis, and limited number of sample points. We also notice that when the low-end of the x-axis for the scatter plot is close to the origin (e.g. drosophila), the slope is relatively larger and closer to 1. Another observation is the difference of vocabulary sizes among different genomes, with mouse having more protein domain types than chicken, zebrafish and drosophila. Word size differences between genomes were also observed in (Gatherer, 2007), though for different definition of words (over-represented protein substrings).
Figure 6: Token-type plots for mouse, zebrafish, chicken, and drosophila genomes. Each black dot represents a chromosome and blue dot for half-chromosome. The chromosome number of some outliers are indicated. The regression slopes from log-log plot are the Heaps’ law exponent.
5.2 Random sampling from total CRCPD pool
Similar to subsection 4.2, we may create our own sampling units by randomly sampling CRCPD tokens from a genome. This step is even more important for other animal genomes than the human genome, because the type-token plots in Fig.6 are more noisy and are more limited in x-axis range.
Figure 7 shows the number of CRCPD types within a randomly sampled CRCPD tokens (without replacement), as a function of number of tokens, for the four genomes (mouse, drosophila, chicken, and zebrafish). The trend of these type-token plots is very similar: quadratic regression lines fit the data perfectly, and within limited range, these quadratic lines can be approximated as linear segments (thus Heaps’ law).
The quadratic nature of the type-token plot when x-axis is extended reveals a weakness of using linear regression to fit the data in Fig.6. The slight difference between the linear coefficients in the different genomes in Fig.6 is not a reliable index to rank genomes. Figure 7, on the other hand, shows a clear order, from zebrafish, to chicken and drosophila, then to mouse, following the trend of increase in vocabulary sizes (number of distinct CRCPD types). The human data is also shown in Fig.7 which is essentially identical to that of mouse.
Figure 7: For the four animal genomes (mouse, drosophila, chicken, and zebrafish), number of unique CRCPD types as a function of the number of randomly sampled CRCPD tokens (in log-log scale). The quadratic regression lines (for log-x and log-y) are also shown.
@@ 6 Discussion
Since the genome size is limited, the total number of protein-domain-coding loci is limited, therefore there is an upper limit in the #token axis as well as in the #type axis. Similarly, due to the large proportion ($\sim$ 50%) of human genome being in non-coding intergenic regions without any CRCPDs, it is not practical to use a short genome region as it may have zero count of #token and #type. This leads to a lower limit in #token axis. Checking type-token relation in the human genome has to be range limited. On the other hand, our simulation in Fig.4 shows that beyond certain range, one should not expect power-law relationship between type and token anymore, and Heaps’ law will break down.
@@
Studies along the same lines with the present investigation include (Nasir et al., 2017; Caetano-Anollés, 2021). These works focus on phylogenetics of different species, and the word used is protein fold superfamilies (FSF). FSF information was provided by the SCOP (structural classification of proteins) database (Murzin et al., 1995; Andreeva et al., 2020). Even if SCOP domain superfamilies and Pfam domains are somewhat similar, the sample points in a type-token plot in (Nasir et al., 2017; Caetano-Anollés, 2021) are different from those in the current study. In (Nasir et al., 2017; Caetano-Anollés, 2021), each point represents a species whereas in our study a point is a chromosome or chromosome arm or a unit smaller than the whole genome. When all species are represented in one plot in (Nasir et al., 2017; Caetano-Anollés, 2021), including viruses, archaea, bacteria, eukaryotes, the scattering of the points (genomes) in log-log scale does not follow one single straight line, but a curve. Instead of treating the curve as consisting of multiple piecewise linear regimes (Nasir et al., 2017; Caetano-Anollés, 2021), another possibility is to use a quadratic fitting function; admittedly, this does not provide an explanation for the reason of deviation from the power-law trend, and it ignores some fundamental differences in genome evolution in different domains.
One key lesson from the Heaps’ law that the rate of new knowledge gain does not linearly increase with the effort, is very relevant to pan-genome studies (Tellelin et al., 2008) – the number of new genes discovered does not increase linearly with the number of genome sequenced. A different kind of diminishing return, concerning the alignment of next-generation-sequencing reads when the read length is increased, is discussed in (Li et al., 2014; Stephens and Iyre, 2018).
Besides the above work on protein fold superfamilies in pan-genomes, Heaps’ law, unlike Zipf’s law (Sheinman et al., 2016), rarely appears in the literature concerning genomics. In (Gatherer, 2007), when over-represented peptide substrings (over-represented k-mers with k in a range of values) are considered as words, linear regression is used to fit the type-token scatter plots. No attempt was made to use the power-law function in (Gatherer, 2007). In (Mukhopadhyay et al., 2006), the fixed-length (k=3) codons are considered as words. This definition of words would easily reach the vocabulary limit when the text length is increased. Therefore, it is not surprising that a different type-token relation is expected.
@@
We have confirmed the result of some previous works that if tokens are randomly sampled from a dictionary (type pool), while type frequency (token per type) follows the Zifp’s law or some modification of the Zipf’s law, the resulting type-token plots, though may have a look of power-law/Heaps’ law, are actually curved (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015). By simulation, we also numerically determined the impact of finite dictionary size, by removing the rare CRCPD types from consideration: the type-token plot becomes more curved with a reduction of dictionary size. The previous attempts to derive the type-token relation, with finite or infinite dictionary size do not present the result in closed-form analytic expression (Eliazar, 2011; Lü et al., 2013; Tunnicliffe and Hunter, 2022). Our quadratic regression provides a simple formula for an empirical description of the systematic deviation from the Heaps’ law.
@@
The definition of “type” or “dictionary word”, or in our case, Pfam domains, is supposed to be given by the database, and synonyms are not merged. For example, there are 13 protein domain names containing the prefix of zf-C2H2, including the top ranking zf-C2H2 (5493 appearances in the genome) and second ranking zf-C2H2_6 (1609 appearance in the genome). The reason of not combining potential synonyms is that it should be a feature of a “dictionary”, and actually, Heaps’ law partially reflects this lexical diversity. If one would combine CRCPD that are “similar” (up to certain degree), the type-token plot and Heaps’ law in human languages should also be modified accordingly.
@@
As argued in (Lynch and Conery, 2003), a consequence of smaller population size is that selection pressure is much less effective in removing individuals without a perfect fit for the given environment. As nonadaptive forces dominated, genomes became less compact, with more redundancies, and inefficiencies were tolerated. On the other hand, this near-neutral random walk in Wright’s landscape allows a species to reach a new adaptive peak, potentially higher than the previous one (Wagner, 2019). Microorganisms have populations several orders of magnitude larger than animals. The higher level of genome redundancy and complexity in animals, versus the lower level of genome complexity in microorganisms, is then expected. For a genome without redundancy, the number of token is roughly equal to the number of types, and Heaps exponent is one. With redundancy, exponent is expected to be less than 1.
It is suggested in (Caetano-Anollés et al., 2017; Caetano-Anollés, 2021) that viral, archaeal and bacterial proteomics have a higher level of vocabulary growth than eukarya, as indicated by the higher value of $\alpha$ exponents (around 0.8). The $\alpha$ exponent for eukarya genomes in (Caetano-Anollés, 2021) is only around 0.1, in contrast with our exponent $\alpha\sim 0.7$ within one genome. It is tempting to use the slope in linear regression to rank different animal genomes in their level of duplication of protein domains. However, the limited x-range makes the estimation of slope less reliable. Also, when x-range is expanded, the linear (i.e., power-law when x and y are not in log-scale) trend becomes quadratic, and there are three parameters to characterize the type-token function, not just one. The similarity of all three fitting parameters for the animal genomes in Fig.7 make them more similar than different.
@@
Despite the deviation from a perfect power-law for type-token plot, or in an alternative wording, that the power-law Heaps’ law is only technically correct in a limited range, there are other important lessons in this study from the linguistic perspective. Just like human languages with hundreds of new English words being created annually, albeit some of them are synonyms to the existing words, new protein domains have been generated during evolution. Even if the new protein domains may not be completely new, being at various degrees similar to the old ones, these may not be that different from the new synonymous words in human languages. Some of the seemingly non-novel new protein domains may continue to evolve and eventually lead to a new meaning and a new structure.